OpenStudy (eiwoh):

A bicycle manufacturing company makes a particular type of bike. Each child bike requires 4 hours to build and 4 hours to test. Each adult bike requires 6 hours to build and 4 hours to test. With the number of workers, the company is able to have up to 120 hours of building time and 100 hours of testing time for a week. If c represents child bikes and a represents adult bikes, determine which system of inequality best explains whether the company can build 10 child bikes and 12 adult bikes in the week.

OpenStudy (eiwoh):

No, because the bike order does not meet the restrictions of 4c + 6a ≤ 120 and 4c + 4a ≤ 100 No, because the bike order does not meet the restrictions of 4c + 4a ≤ 120 and 6c + 4a ≤ 100 Yes, because the bike order meets the restrictions of 4c + 6a ≤ 120 and 4c + 4a ≤ 100 Yes, because the bike order meets the restrictions of 4c + 4a ≤ 120 and 6c + 4a ≤ 100

OpenStudy (eiwoh):

@brooke..help00 @KJSaif

OpenStudy (eiwoh):

@mhchen

OpenStudy (brooke..help00):

What do you think?

OpenStudy (eiwoh):

This is a test i have taken a bunch of times but always failed :( but i think it was C

OpenStudy (mhchen):

So each child bike is represented by 'c', each adult bike is represented by 'a'. This is about time so.. 4c = building time of child bike 6a = building time of adult bike 120 = total building time allowed 4c + 6a = 120 4c = testing time of child bike 4c = testing time of adult bike 100 = total testing time allowed 4c + 4a = 100 Now it asks for whether 10 child bikes and 12 adult bikes can be built in a week. In order to test this, you have to make it fit BOTH equations. So we test them: First equation: 4(10) + 6(12) <= 120? Second equation: 4(10) + 6(10) <= 100? I put <= because they can be lazy and produce less, but they can't produce more ofc.

OpenStudy (mhchen):

So First equation: 40 + 80 <= 120? <-- yes Second equation: 40 + 60 <= 100? <-- yes So then you look at your answer choices, and find the one that matches our 2 equations.

OpenStudy (eiwoh):

So C then :)

OpenStudy (mhchen):

ye